# VGLD: Visually-Guided Linguistic Disambiguation for Monocular Depth Scale Recovery

### v2

@@ 8 Zero-shot Generalization
ZoeDepth[5] robust depth $\dagger$ 0.123 0.356 0.053 0.856 0.979 0.995
ScaleDepth[64] 0.129 0.359 0.866
MiDas-1[6] Least Squares 0.197 0.418 0.346 0.278 0.061 0.873 0.964 0.981
Levenberg Marquardt 0.158 0.440 0.252 0.116 0.032 0.950 0.988 0.995
[…]
VGLD-NK-T(Ours) 0.318 0.647 0.566 0.242 0.089 0.643 0.914 0.980
VGLD-NK-I(Ours) 0.259 0.595 0.468 0.202 0.071 0.751 0.957 0.991
VGLD-NK-TCI(Ours) 0.262 0.628 0.467 0.202 0.071 0.751 0.959 0.991
MiDas-2[39] Least Squares 0.203 0.419 0.365 0.272 0.062 0.860 0.961 0.981
Levenberg Marquardt 0.173 0.438 0.291 0.132 0.039 0.926 0.984 0.994
[…]
VGLD-NK-I(Ours) 0.288 0.688 0.552 0.246 0.090 0.627 0.922 0.984
VGLD-NK-TCI(Ours) 0.275 0.670 0.513 0.225 0.080 0.694 0.941 0.987
DAV2-vits[54] Least Squares 0.194 0.418 0.337 0.305 0.062 0.880 0.963 0.980
Levenberg Marquardt 0.146 0.439 0.224 0.103 0.027 0.961 0.989 0.995
[…]
VGLD-NK-I(Ours) 0.273 0.564 0.535 0.236 0.090 0.617 0.931 0.989
VGLD-NK-TCI(Ours) 0.241 0.545 0.433 0.189 0.067 0.779 0.967 0.993
DAV1-vits[53] Least Squares 0.196 0.416 0.341 0.282 0.061 0.875 0.963 0.981
Levenberg Marquardt 0.151 0.440 0.234 0.108 0.029 0.957 0.989 0.995
[…]
VGLD-NK-T(Ours) 0.281 0.583 0.532 0.214 0.078 0.711 0.945 0.987
VGLD-NK-I(Ours) 0.250 0.573 0.443 0.194 0.070 0.764 0.965 0.991
VGLD-NK-TCI(Ours) 0.241 0.545 0.433 0.189 0.067 0.779 0.967 0.993
Table 7: Zero-shot generalization to SUN-RGBD (Indoor). The yellow regions represent the results obtained using least-squares-type mathematical fitting methods. Best results are in bold, second best are underlined.
RDE Model Method Lower is better Higher is better
MiDas-1[6] Least Squares 0.123 0.070 0.364 0.357 0.069 0.868 0.959 0.978
Levenberg Marquardt 0.070 0.029 0.241 0.095 0.029 0.952 0.991 0.998
[…]
VGLD-NK-T (Ours) 0.251 0.385 0.683 0.240 0.094 0.622 0.898 0.969
VGLD-NK-I (Ours) 0.188 0.138 0.544 0.208 0.079 0.696 0.943 0.982
VGLD-NK-TCI (Ours) 0.212 0.281 0.623 0.228 0.088 0.638 0.930 0.978
MiDas-2[39] Least Squares 0.133 0.080 0.394 0.345 0.071 0.846 0.954 0.977
Levenberg Marquardt 0.086 0.039 0.285 0.114 0.036 0.929 0.988 0.996
[…]
VGLD-NK-I (Ours) 0.235 0.201 0.722 0.294 0.115 0.460 0.849 0.975
VGLD-NK-TCI (Ours) 0.227 0.371 0.690 0.262 0.100 0.570 0.894 0.979
DAV2-vits[54] Least Squares 0.123 0.068 0.361 0.361 0.069 0.872 0.960 0.978
Levenberg Marquardt 0.066 0.024 0.226 0.092 0.028 0.958 0.993 0.998
[…]
VGLD-NK-I (Ours) 0.212 0.169 0.663 0.259 0.103 0.514 0.899 0.989
VGLD-NK-TCI (Ours) 0.196 0.487 0.610 0.208 0.082 0.678 0.952 0.990
DAV1-vits[53] Least Squares 0.118 0.063 0.345 0.344 0.066 0.875 0.961 0.979
Levenberg Marquardt 0.056 0.020 0.203 0.081 0.024 0.970 0.994 0.999
[…]
VGLD-NK-T (Ours) 0.211 0.711 0.627 0.215 0.084 0.683 0.927 0.983
VGLD-NK-I (Ours) 0.193 0.200 0.597 0.220 0.087 0.619 0.950 0.994
VGLD-NK-TCI (Ours) 0.196 0.487 0.610 0.208 0.082 0.678 0.952 0.990
Table 8: Zero-shot generalization to DIML Indoor. The yellow regions represent the results obtained using least-squares-type mathematical fitting methods. Best results are in bold, second best are underlined.
RDE Model Method Lower is better Higher is better
MiDas-1[6] Least Squares 0.123 0.070 0.364 0.357 0.069 0.868 0.959 0.978
Levenberg Marquardt 0.070 0.029 0.241 0.095 0.029 0.952 0.991 0.998
[…]
VGLD-NK-T (Ours) 0.251 0.385 0.683 0.240 0.094 0.622 0.898 0.969
VGLD-NK-I (Ours) 0.188 0.138 0.544 0.208 0.079 0.696 0.943 0.982
VGLD-NK-TCI (Ours) 0.212 0.281 0.623 0.228 0.088 0.638 0.930 0.978
MiDas-2[39] Least Squares 0.133 0.080 0.394 0.345 0.071 0.846 0.954 0.977
Levenberg Marquardt 0.086 0.039 0.285 0.114 0.036 0.929 0.988 0.996
[…]
VGLD-NK-I (Ours) 0.235 0.201 0.722 0.294 0.115 0.460 0.849 0.975
VGLD-NK-TCI (Ours) 0.227 0.371 0.690 0.262 0.100 0.570 0.894 0.979
DAV2-vits[54] Least Squares 0.123 0.068 0.361 0.361 0.069 0.872 0.960 0.978
Levenberg Marquardt 0.066 0.024 0.226 0.092 0.028 0.958 0.993 0.998
[…]
VGLD-NK-I (Ours) 0.212 0.169 0.663 0.259 0.103 0.514 0.899 0.989
VGLD-NK-TCI (Ours) 0.196 0.487 0.610 0.208 0.082 0.678 0.952 0.990
DAV1-vits[53] Least Squares 0.118 0.063 0.345 0.344 0.066 0.875 0.961 0.979
Levenberg Marquardt 0.056 0.020 0.203 0.081 0.024 0.970 0.994 0.999
[…]
VGLD-NK-T (Ours) 0.211 0.711 0.627 0.215 0.084 0.683 0.927 0.983
VGLD-NK-I (Ours) 0.193 0.200 0.597 0.220 0.087 0.619 0.950 0.994
VGLD-NK-TCI (Ours) 0.196 0.487 0.610 0.208 0.082 0.678 0.952 0.990
Table 9: Zero-shot generalization to DDAD (Outdoor). The yellow regions represent the results obtained using least-squares-type mathematical fitting methods.
9 Ablation Study
9.1 Effect of the DRM
As shown in Table10, we conduct ablation studies on the Domain Router Mechanism (DRM). The results demonstrate that the inclusion of DRM consistently improves the overall performance of VGLD and significantly enhances its cross-domain generalization capabilities.
RDE Model Method NYU KITTI
[…]
MiDas-1[6] w/o DRM 0.128 0.415 0.757 0.13 3.636 0.855
with DRM 0.120 0.414 0.863 0.12 3.559 0.874
MiDas-2[39] w/o DRM 0.158 0.513 0.740 0.19 4.987 0.728
with DRM 0.151 0.513 0.780 0.18 4.804 0.737
DAV2-vits[54] w/o DRM 0.135 0.459 0.798 0.163 4.060 0.767
with DRM 0.127 0.434 0.835 0.153 3.980 0.772
DAV1-vits[53] w/o DRM 0.122 0.437 0.847 0.145 4.327 0.748
with DRM 0.112 0.392 0.883 0.136 4.008 0.816
Table 10: Performance comparison on NYU and KITTI datasets with and w/o DRM. Best results are in bold.
9.2 Effect of the LM loss
To investigate the effect of different weights of LM loss $\mathcal{L}_{lm}$ on model training, we vary the value of $\alpha$ and train the VGLD-NK-TCI model based on the DAV1-vits RDE backbone. Evaluation on both the NYUv2 and KITTI datasets shown in Table11 that the model achieves the best performance when $\alpha=0.1$.
$\alpha$ NYU KITTI
[…]
0 0.115 0.403 0.874 0.164 4.856 0.781
0.001 0.116 0.413 0.869 0.161 4.204 0.779
0.01 0.113 0.399 0.879 0.146 4.010 0.791
0.1 0.112 0.392 0.883 0.136 4.008 0.816
1 0.115 0.397 0.868 0.162 4.701 0.778
9.3 Effect of the initial seeds.
@@ 9.2 Effect of the LM loss
10 Computational Complexity
As shown in Table 12, we present a comparison of model parameters and inference time between VGLD and RSA to quantify the computational resources required. All evaluations were conducted on a single NVIDIA RTX 3090 (24GB). The results indicate that the scalar predictor in VGLD is more lightweight than RSA. However, VGLD additionally incorporates a CLIP image encoder, which introduces an extra 14ms of inference time compared to RSA. Despite this overhead, VGLD offers a favorable trade-off: it achieves a 32.1%(Ref. to Tabel 1) improvement in Abs Rel on NYUv2 with an inference time of just 39ms and a modest parameter, making it a practical and efficient choice.
Components RSA[59] VGLD (ours)
Params# Inference Time Params# Inference Time
DAV1-vits[53] 24.78M 9.62ms 24.78M 9.62ms
CLIP Text Encoder[38] 63.43M 13.61ms 63.43M 13.61ms
CLIP Image Encoder[38] - - 86.19M 14.90ms
Scalars Predictor 1.49M 1.76ms 1.18M 0.94ms
Total 89.7M 24.99ms 175.58M 39.07ms
Increase / M (ms) - - 85.88M $\uparrow$ 14.08ms $\uparrow$
Table 12: Computational Complexity Analysis. As shown in the table, the increase in model parameters and inference time of VGLD compared to the RSA model primarily stems from the additional CLIP Image Encoder component.
11 Prompts for Natural Text Generation
To generate natural and semantically rich image descriptions—rather than relying on fixed prompt templates—we employ two vision-language models: LLaVA-v1.6-Vicuna-7B and LLaVA-v1.6-Mistral-7B[20]. To ensure diversity in the generated captions, each model is prompted using six distinct instruction templates. These prompt templates are listed in Table 13.
@@ 11 Prompts for Natural Text Generation
Table 13: Prompts for Natural Text Generation. We utilize two LLaVA models, each generating 6 textual descriptions per image, resulting in a total of 12 diverse descriptions for each image.

### v3

@@ Appendix A Supplementary Material
RDE Model Method Lower is better Higher is better
MiDas-1(Reiner et al. 2023) Least Squares 0.123 0.070 0.364 0.357 0.069 0.868 0.959 0.978
Levenberg Marquardt 0.070 0.029 0.241 0.095 0.029 0.952 0.991 0.998
RSA-NK(Zeng et al. 2024b) 0.219 0.218 0.667 0.246 0.096 0.612 0.882 0.964
VGLD-NK-T (Ours) 0.251 0.385 0.683 0.240 0.094 0.622 0.898 0.969
VGLD-NK-I (Ours) 0.188 0.138 0.544 0.208 0.079 0.696 0.943 0.982
VGLD-NK-TCI (Ours) 0.212 0.281 0.623 0.228 0.088 0.638 0.930 0.978
MiDas-2(Ranftl et al. 2020) Least Squares 0.133 0.080 0.394 0.345 0.071 0.846 0.954 0.977
Levenberg Marquardt 0.086 0.039 0.285 0.114 0.036 0.929 0.988 0.996
VGLD-NK-T (Ours) 0.243 0.359 0.737 0.264 0.100 0.585 0.877 0.964
VGLD-NK-I (Ours) 0.235 0.201 0.722 0.294 0.115 0.460 0.849 0.975
VGLD-NK-TCI (Ours) 0.227 0.371 0.690 0.262 0.100 0.570 0.894 0.979
DAV2-vits(Yang et al. 2024b) Least Squares 0.123 0.068 0.361 0.361 0.069 0.872 0.960 0.978
Levenberg Marquardt 0.066 0.024 0.226 0.092 0.028 0.958 0.993 0.998
VGLD-NK-T (Ours) 0.228 0.300 0.673 0.246 0.096 0.593 0.891 0.981
VGLD-NK-I (Ours) 0.212 0.169 0.663 0.259 0.103 0.514 0.899 0.989
VGLD-NK-TCI (Ours) 0.196 0.487 0.610 0.208 0.082 0.678 0.952 0.990
DAV1-vits(Yang et al. 2024a) Least Squares 0.118 0.063 0.345 0.344 0.066 0.875 0.961 0.979
Levenberg Marquardt 0.056 0.020 0.203 0.081 0.024 0.970 0.994 0.999
RSA-NK(Zeng et al. 2024b) 0.216 0.283 0.679 0.249 0.098 0.608 0.873 0.964
VGLD-NK-T (Ours) 0.211 0.711 0.627 0.215 0.084 0.683 0.927 0.983
VGLD-NK-I (Ours) 0.193 0.200 0.597 0.220 0.087 0.619 0.950 0.994
VGLD-NK-TCI (Ours) 0.196 0.487 0.610 0.208 0.082 0.678 0.952 0.990
Table 12: Zero-shot generalization to DIML Indoor. Best results are in bold, second best are underlined.
@@
MiDas-1(Reiner et al. 2023) Least Squares 0.319 2.265 7.252 1.936 0.301 0.409 0.844 0.920
Levenberg Marquardt 0.201 1.231 5.411 0.223 0.079 0.673 0.960 0.991
RSA-NK(Zeng et al. 2024b) 0.223 - 19.342 0.325 - 0.631 0.903 0.966
VGLD-NK-T (Ours) 0.215 2.519 10.467 0.320 0.102 0.630 0.851 0.935
VGLD-NK-I (Ours) 0.212 2.409 10.061 0.311 0.101 0.633 0.851 0.935
VGLD-NK-TCI (Ours) 0.209 2.517 10.446 0.319 0.100 0.659 0.862 0.941
MiDas-2(Ranftl et al. 2020) Least Squares 0.328 2.447 7.490 1.902 0.298 0.407 0.828 0.914
Levenberg Marquardt 0.232 1.557 5.985 0.253 0.090 0.609 0.934 0.985
VGLD-NK-T (Ours) 0.232 2.625 14.324 0.326 0.112 0.603 0.841 0.936
VGLD-NK-I (Ours) 0.220 2.526 12.235 0.321 0.106 0.642 0.865 0.947
VGLD-NK-TCI (Ours) 0.212 2.521 10.032 0.311 0.102 0.659 0.881 0.954
DAV2-vits(Yang et al. 2024b) Least Squares 0.318 2.239 7.205 1.937 0.300 0.410 0.847 0.920
Levenberg Marquardt 0.173 1.027 4.988 0.200 0.069 0.757 0.974 0.992
VGLD-NK-T (Ours) 0.221 3.125 8.769 0.252 0.085 0.675 0.927 0.977
VGLD-NK-I (Ours) 0.185 2.848 8.344 0.232 0.074 0.746 0.929 0.980
VGLD-NK-TCI (Ours) 0.176 2.002 7.925 0.238 0.075 0.748 0.942 0.981
DAV1-vits(Yang et al. 2024a) Least Squares 0.316 2.223 7.182 1.932 0.299 0.411 0.850 0.920
Levenberg Marquardt 0.156 0.929 4.766 0.185 0.062 0.817 0.977 0.991
RSA-NK(Zeng et al. 2024b) 0.207 - 19.715 0.303 - 0.642 0.903 0.976
VGLD-NK-T (Ours) 0.210 2.598 13.432 0.318 0.108 0.708 0.913 0.970
VGLD-NK-I (Ours) 0.192 2.557 9.275 0.258 0.081 0.732 0.922 0.975
VGLD-NK-TCI (Ours) 0.186 2.403 8.984 0.246 0.079 0.742 0.932 0.975
Table 13: Zero-shot generalization to DDAD (Outdoor). Best results are in bold, second best are underlined.
@@
Table 14: Prompts for Natural Text Generation. We utilize two LLaVA models(llava-v1.6-vicuna-7b and llava-v1.6-mistral-7b), each generating 6 textual descriptions per image, resulting in a total of 12 diverse descriptions for each image.